# Scalable Verification of Neural Control Barrier Functions Using Linear Bound Propagation

### v1

@@ 1 Introduction
To overcome these challenges, CBFs represented as neural networks—neural CBFs—have become a popular alternative, enabling the inductive synthesis of barrier functions for a broad class of dynamics and safety constraints (Zhao et al., 2020; Zhang et al., 2023; Hu et al., 2024; Abate et al., 2021b). Neural CBFs allow for arbitrary nonlinear dynamics, and the expressiveness of the barrier function can be increased with the neural network’s size, thereby expanding the obtainable control invariant set. However, the main drawback of learning a candidate barrier function arguably lies in the high cost of post-hoc verification required to prove that the candidate is a valid CBF. This verification stage constitutes the primary computational bottleneck, constrained by two key factors: (1) the size of the neural network and (2) the dimensionality of the dynamical system. Existing verification approaches predominantly rely on satisfiability modulo theories (SMT) solvers (Zhao et al., 2020; Abate et al., 2021a; Sha et al., 2021; Edwards et al., 2024) and mixed-integer programming (MIP) (Zhao et al., 2022), both of which exhibit poor scalability with respect to network size. Architecture-specific approaches (Zhang et al., 2023; Hu et al., 2024), meanwhile, lack generalization to networks with arbitrary nonlinear activation functions needed for sufficient expressivity.
@@ 4.2 Linear Bounds on the CBF Constraint
$\frac{\partial\mathcal{B}_{\theta}(x)}{\partial x}f(x)$ $\geq\Gamma(\eta)\,x+\beta(\eta),$ (10)
@@
$\psi_{\mathrm{invar}}(x)\coloneqq\left[\Gamma_{\mathrm{drift,L}}(\eta)+\Gamma_{\mathrm{ctrl,L}}(\eta)+\alpha\underline{\Pi}\,\right]\,x+\beta_{\mathrm{drift,L}}(\eta)+\beta_{\mathrm{ctrl,L}}(\eta)+\alpha\underline{\pi}\geq 0.$ (15)
@@
$\psi_{\mathrm{invar}}^{\mathrm{NEG}}(x)\coloneqq\left[\Gamma_{\mathrm{drift,U}}(\eta)+\Gamma_{\mathrm{ctrl,U}}(\eta)+\alpha\overline{\Pi}\,\right]x+\beta_{\mathrm{drift,U}}(\eta)+\beta_{\mathrm{ctrl,U}}(\eta)+\alpha\overline{\pi}<0.$ (17)
@@ 4.3 Refinement of the Domains for Linear Bounds
Our approach enables efficient utilization of computational resources by performing LBP on batches of simplices using GPUs. In particular, we collect simplices that require evaluation of $\psi_{\mathrm{invar}}$ into batches and compute the necessary linear bounds on the invariance condition jointly over each batch. By further parallelizing verification over simplices, potentially over multiple GPUs, our approach allows for efficient utilization of computational resources to handle more challenging verification tasks involving larger networks.
@@ 5 Experimental Results
We evaluate the effectiveness of our proposed approach in scaling to larger neural network architectures. We adopt standard benchmarks from the literature (Zeng et al., 2016; Zhang et al., 2023; Abate et al., 2021a; Liu et al., 2015; Prajna, 2006; Zhao et al., 2020; Barry et al., 2012), along with two novel benchmarks (2D Control and Cart Pole) chosen to demonstrate the incorporation of the $\sup_{u}g(x)u$ term in Eq. 2. Additional details on all benchmarks are provided in Appendix E. Across all experiments, we employ the tanh activation function, as it consistently yields the largest control invariant sets. Implementations of the corresponding relaxations for the sigmoid, ReLU, and leaky-ReLU activation functions are included in the accompanying codebase [^1] The code can be found at https://github.com/Zinoex/verification-of-neural-cbf-via-lbp.. We fix the McCormick relaxation parameter to $\eta=\tfrac{1}{2}$, and class-$\mathcal{K}$ function $\alpha=1$ for all experiments. All computations ran on a machine with an Intel i7-6700K CPU, 16 GB RAM, and an NVIDIA GTX 1060 GPU (6 GB VRAM). Details regarding network training procedures are provided in Appendix A.
@@ Results
Figure 2: Illustration of the neural control barrier function $\mathcal{B}_{\theta}(x)$, the CBF invariance condition from Eq. 2, and the resulting verification regions after adaptive mesh refinement.
Fig. 2 presents the verification results for the Darboux benchmark. As shown in this figure, regions where the invariance condition (Eq. 15) or the value of $\mathcal{B}_{\theta}$ are close to zero require the most refinement, while the mesh can remain coarse in regions where even conservative linearization is adequate to establish validity. The figures for the other benchmarks are presented in Appendix E.
@@ 6 Conclusion
We have presented a scalable verification framework for neural control barrier functions that effectively handles larger neural network architectures and incorporates control inputs into the verification process. By leveraging linear bound propagation and first-order Taylor expansions, we constructed upper and lower linear bounds on the nonlinearity inherent in the verification task, thereby circumventing the need for SMT solvers capable of reasoning over quantifier-free nonlinear real arithmetic formulae. Our numerical experiments on a variety of benchmarks have demonstrated that our approach applies to a wide class of dynamics and to nonlinear activation functions, outperforms existing SMT-based verification tools, and scales to significantly larger networks. Future work will consider extending our verification framework to certificates of specifications beyond invariance, such as for reachability and avoidance.

### v2

@@ 1 Introduction
To overcome these challenges, CBFs represented as neural networks—neural CBFs—have become a popular alternative, enabling the inductive synthesis of barrier functions for a broad class of dynamics and safety constraints (Zhao et al., 2020; Zhang et al., 2023; Hu et al., 2024; Abate et al., 2021b). Neural CBFs allow for arbitrary nonlinear dynamics, and the expressiveness of the barrier function can be increased with the neural network’s size, thereby expanding the obtainable control invariant set. However, the main drawback of learning a candidate barrier function arguably lies in the high cost of post-hoc verification required to prove that the candidate is a valid CBF. This verification stage constitutes the primary computational bottleneck, constrained by two key factors: (1) the size of the neural network and (2) the dimensionality of the dynamical system. Existing verification approaches that rely on satisfiability modulo theories (SMT) solvers (Zhao et al., 2020; Abate et al., 2021a; Sha et al., 2021; Edwards et al., 2024) or mixed-integer programming (MIP) (Zhao et al., 2022) exhibit poor scalability with respect to network size. Approaches that have employed branch-and-bound verification have until now been restricted to ReLU-based neural CBFs, as this facilitates efficient computation of the gradient of the network (Shi et al., 2022; Zhang et al., 2023; Hu et al., 2024). Learning neural CBFs for challenging problems requires introducing sufficient nonlineararity into the network, which motivates our focus on networks with arbitrary nonlinear activation functions.
@@ 4.2 Linear Bounds on the CBF Constraint
$\psi_{\mathrm{invar}}(x)\coloneqq\left[\Gamma_{\mathrm{drift,L}}(\eta)+\Gamma_{\mathrm{ctrl,L}}(\eta)+\alpha\underline{A}_{\mathcal{B}_{\theta}}\,\right]x+\beta_{\mathrm{drift,L}}(\eta)+\beta_{\mathrm{ctrl,L}}(\eta)+\alpha\underline{a}_{\mathcal{B}_{\theta}}\geq 0.$ (15)
@@
$\psi_{\mathrm{invar}}^{\mathrm{NEG}}(x)\coloneqq\left[\Gamma_{\mathrm{drift,U}}(\eta)+\Gamma_{\mathrm{ctrl,U}}(\eta)+\alpha\overline{A}_{\mathcal{B}_{\theta}}\,\right]x+\beta_{\mathrm{drift,U}}(\eta)+\beta_{\mathrm{ctrl,U}}(\eta)+\alpha\overline{a}_{\mathcal{B}_{\theta}}<0.$ (17)
@@ 4.3 Refinement of the Domains for Linear Bounds
Our approach enables efficient utilization of computational resources by performing LBP on batches of simplices using GPUs. In particular, we collect simplices that require evaluation of the SAT formula $\psi_{\mathrm{invar}}$ defined in Eq. 15 into batches and compute the necessary linear bounds on the invariance condition jointly over each batch. By further parallelizing verification over simplices, potentially over multiple GPUs, our approach allows for efficient utilization of computational resources to handle more challenging verification tasks involving larger networks.
@@ 5 Experimental Results
We evaluate the effectiveness of our proposed approach in scaling to larger neural network architectures. We adopt standard benchmarks from the literature (Zeng et al., 2016; Zhang et al., 2023; Abate et al., 2021a; Liu et al., 2015; Prajna, 2006; Zhao et al., 2020; Barry et al., 2012), along with two novel benchmarks (2D Control and Cart Pole) chosen to demonstrate the incorporation of the $\sup_{u}g(x)u$ term in Eq. 2. Additional details on all benchmarks are provided in Appendix E. Across all experiments, we employ the tanh activation function, as it consistently yields the largest control invariant sets. Implementations of the corresponding relaxations for the sigmoid, ReLU, and leaky-ReLU activation functions are included in the accompanying codebase [^1] The code can be found at https://github.com/Zinoex/verification-of-neural-cbf-via-lbp.. We fix the McCormick relaxation parameter to $\eta=\tfrac{1}{2}$, and class-$\mathcal{K}$ function $\alpha=1$ for all experiments. One could choose to optimize $\eta$, e.g., using projected gradient descent, similar to the optimization of parameterized linear relaxations in $\alpha$-CROWN (Xu et al., 2020), though for simplicity we fix it to a constant. All computations ran on a machine with an Intel i7-6700K CPU, 16 GB RAM, and an NVIDIA GTX 1060 GPU (6 GB VRAM). Details regarding network training procedures are provided in Appendix A.
@@ Results
Figs. 8, 9, 10 and 11 present the verification results for the benchmarks. As shown in these figures, regions of the state space where the invariance condition (Eq. 15) or the value of $\mathcal{B}_{\theta}$ are close to zero require the most refinement, while the mesh can remain coarse in regions where even conservative linearization is sufficient to establish the validity of the neural CBF.
@@ 6 Conclusion
We have presented a scalable verification framework for neural control barrier functions that effectively handles larger neural network architectures and incorporates control inputs into the verification process. By leveraging linear bound propagation and first-order Taylor expansions, we constructed upper and lower linear bounds on the nonlinearity inherent in the verification task, thereby circumventing the need for SMT solvers capable of reasoning over quantifier-free nonlinear real arithmetic formulae. Our numerical experiments on a variety of benchmarks have demonstrated that our approach applies to a wide class of dynamics and to nonlinear activation functions, outperforms existing SMT-based verification tools, and scales to significantly larger neural networks. Future work will consider extending our verification framework to certificates of specifications beyond invariance, such as for reachability and avoidance.